# Language Models That Walk the Talk: A Framework for Formal Fairness Certificates

### v1

@@ 1 Introduction
As large language models become integral to academia and industry, particularly in high-stakes decision-making and safety-critical domains, ensuring their robustness, alignment, and trustworthiness is crucial garcia2024trustworthy ; huang2024survey . While their transformer-based architecture with attention mechanisms has driven widespread adoption vaswani2017attention , concerns about robustness and safety persist, especially in tasks like fairness enforcement and content moderation.
A key challenge is their susceptibility to adversarial perturbations goodfellow2015explaining ; szegedy2013intriguing , where small perturbations, such as synonym substitutions or paraphrasing, can manipulate model behavior alzantot2018generating ; behjati2019universal ; li2019textbugger ; liang2017deep ; jin2020bert ; hsieh2019robustness . While such perturbations may be adversarially crafted in settings like toxicity detection bespalov2023towards ; schmidhuber2024llm , they can also arise naturally in fairness-critical tasks such as salary prediction, where gender bias may occur, highlighting the need for robust gender bias mitigation tumlin2024fairnnv . Many natural language processing tasks operate in discrete input spaces (language), where verifying robustness by enumerating all synonym substitutions becomes computationally infeasible due to combinatorial explosions. This holds especially for longer text sequences and words with many synonyms, which makes adversarial robustness verification particularly challenging bonaert2021fast ; huang2019achieving . Specifically, the number of possible 20-token-long adversarial sequences drawn from a 32k-token (BERT-size) vocabulary is equal to $32\text{k}^{20}\approx 10^{90}$, a figure that exceeds the estimated number of atoms in the observable universe kumar2024certifying .
Formal verification brix2023yearsinternationalverificationneural of neural networks has gained significant interest in recent years. These methods rigorously prove that no adversarial input exists within some input set, ensuring that model behavior remains unchanged under perturbations. However, existing approaches have not been extensively applied to large language models and do not provide a holistic view to verify them, particularly in the context of fairness and bias mitigation.
This work provides a holistic verification framework to determine the fairness of transformer-based language models. To summarize our contributions, we first propose a pre-training technique that encourages synonyms, gender-related terms, or toxic terms to reside in close proximity within the embedding space, respectively. Secondly, we integrate the pre-trained embedding layer into a transformer-based classification language model, which is subsequently trained on downstream classification tasks. Thirdly, we assess whether the models maintain their classifications when certain words are substituted with other gender-related terms or adversarially crafted synonyms, using a novel, coverage-based fairness score. This framework is rigorously demonstrated in a fairness-critical task involving gender bias mitigation and a safety-critical task involving toxicity detection.
@@ 2 Background
We consider a general neural network classifier architecture that supports any element-wise nonlinear activation function. Given an input $x\in\mathbb{R}^{n}$, the classifier is represented by $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{c}$. A variety of tools have been developed to formally verify neural network properties, with adversarial robustness being the most extensively studied aspect brix2023yearsinternationalverificationneural . The verification problem for neural networks can be formally stated as follows:
@@
Exact verification of robustness properties in neural networks is often computationally infeasible katz2017reluplex . Therefore, sound but incomplete methods are usually used for larger networks katz2019marabou ; zhang2018efficient ; bak2021nnenum ; lopez2023nnv ; singh2018fast ; duong2024harnessing . In particular, we use zonotopes girard2005reachability ; bonaert2021fast to enclose the output of a neural network in this work, which has been shown to be a suitable foundation for analyzing larger neural networks singh2018fast ; 8418593 ; mirman2018differentiable . As zonotopes offer an efficient and scalable way to represent perturbed input (or embedding) spaces, they have also been used to verify transformer-based models bonaert2021fast . A detailed explanation of zonotope-based neural network verification can be found in Appendix A. However, our work is agnostic to the underlying verification engine and others can be used as well brix2023fourth ; brix2024fifth .
@@ 3 Methodology
This section presents the fairness verification framework employed in this work. As a running example, we illustrate the process using gender fairness verification in a salary prediction task. A visual overview of the framework is provided in Figure 1.
Figure 1: Overview of the framework explained in section 3 involving transfer learning (step 1 and 2), transformer integration & training (step 3), and formal verification (step 4). The transformer architecture in step 3 is from vaswani2017attention .
@@
We begin by pre-training an embedding layer on curated gendered word pairs, encouraging proximity between gendered pairs while maximizing distance from gender-neutral words. This enforces semantic neutrality and fairness at the embedding level through $\ell_{2}$ regularization.
@@
The pre-trained embedding layer is integrated into a transformer-based language model and kept frozen during downstream training of the model, then the language model is trained on a fairness-critical task involving gender bias mitigation and a safety-critical task involving toxicity detection.
@@ 3.1 Pre-training the embedding model
Since many AI-integrated automated decisions (including which individuals will receive jobs, loans, and medication) can significantly impact people’s lives, it is paramount to assess and improve the ethics of the decisions made by these automated systems 10.1145/3494672 ; 10.1145/3457607 . A straightforward approach to promoting fairness in machine learning, defined as the absence of bias and discrimination, is to preclude the use of sensitive attributes during model training and inference lepri2018fair ; calders2010three ; kamiran2010discrimination ; barocas2016big . For instance, in designing a race-blind or gender-blind decision-making system, one might either exclude sensitive attributes, such as gender, from the feature set entirely lepri2018fair , or alternatively, represent them with similar embeddings that capture shared semantic meaning, placing them in the same semantic class to ensure equivalent treatment by the model bolukbasi2016man .
@@
Word embedding models map words to vectors based on the assumption that words appearing in similar contexts share similar meanings 10.1145/3494672 ; mikolov2013efficient ; mikolov2013distributed . Trained on large corpora, they capture semantic relationships and are widely used in natural language processing tasks such as search, translation, and classification 10.1145/3494672 . However, studies have shown that embeddings can encode biases bolukbasi2016man ; brunet2019understanding ; Caliskan_2017 ; zhao2018learning ; for instance, computer programmer may appear closer to male than female, raising concerns about potential discrimination in AI-based decision-making systems.
@@
This work employs pre-training-based transfer learning to train an embedding layer from scratch and cluster gender-related words while distancing unrelated ones. This offers greater transparency, controllability, and verifiability, enabling direct mitigation of gender bias during embedding layer training by mapping all gender-related words to a cohesive region in the embedding space. This semantic alignment enables formal verification. Similarly, this framework can also be applied to adversarial substitutions in the context of toxicity detection, ensuring the model reliably detects and censors inappropriate content. The motivation for this transfer learning design also stems from the need to better capture semantic relationships—especially those involving gender-related, synonymous, or toxic terms—which are not always learned through backpropagation alone [^1] The necessity of employing transfer learning through pre-training arises from the fact that some embedding layers (such as Keras) are randomly initialized at the start of model training keras2023 . Consequently, any semantic understanding is learned solely through backpropagation of transformer model training during gradient updates, which is typically insufficient to capture complex relationships, such as those involving gender or toxicity.. Prior work supports this, showing embeddings trained on task-specific data can outperform static embeddings like Word2Vec 10.1145/3440840.3440847 . Similar practices are seen in domains such as fake news detection liu2018early , highlighting the broader value of transfer learning to improve task performance and propagate semantic knowledge across domains 9134370 . We extend this to the context of toxic language, encompassing a broad range of categories such as profanity, hate speech, abusive expressions, and violent or otherwise harmful language, enabling the model to learn semantics relevant to both bias mitigation and toxicity detection.
@@
The performance of a machine learning system heavily depends on its data representation, with certain feature sets assumed to capture fundamental characteristics effectively across tasks 9226466 . An effective representation should leverage label information by encouraging intra-class compactness, pulling together samples from the same class in the embedding space, while promoting inter-class separability by pushing apart samples from different classes khosla2020supervised ; jaiswal2020survey ; haochen2021provable ; wang2021understanding . Prior research has established contrastive learning as a highly effective component in natural language processing, visual computing, and other domains, demonstrating its ability to enhance representation learning across a wide range of tasks jaiswal2020survey ; caron2020unsupervised ; he2020momentum ; chen2020simple ; haochen2021provable ; saunshi2019theoretical ; chen2020big ; chen2020improved ; he2020momentum ; 1467314 . To learn semantically meaningful word representations, we employ a contrastive loss function designed to differentiate between semantically similar and dissimilar word pairs in the embedding space.
@@
The loss function used in this work is inspired by the formulation introduced in 1467314 and has been adapted to suit our specific task. Here, the predefined margin $m>0$ specifies the minimum required separation between dissimilar pairs, while the hyperparameter $\alpha\geq 0$ regulates the influence of an additional linear attraction term applied to similar pairs. The loss encourages the model to minimize distances between similar pairs while enforcing a separation of at least of at least $m$ units between dissimilar pairs in the embedding space.
Examples of gender-related terms are given in Section B.1. More details on the contrastive loss function can be found in Section B.2. Detailed explanation on the data construction of word pairs for the pre-training phase can be found in Section B.3. All additional embedding pre-training, such as those involving synonym and antonym relationships, was conducted using the Synset structure provided by WordNet 10.1145/219717.219748 , which offers a rich lexical database for modeling semantic similarity and contrast.
@@ 3.2 Training the language model
The pre-trained embedding layer is subsequently integrated into a custom transformer-based language model, inspired by the architecture proposed in vaswani2017attention . Notably, the pre-trained embedding layer is frozen during the transformer training phase and is not updated via backpropagation. This is a common practice, e.g., in models like ELMo peters2018deepcontextualizedwordrepresentations , to preserve language knowledge learned during pre-training and using its representations in downstream tasks. Similarly, ULMFiT howard2018universallanguagemodelfinetuning gradually fine-tunes a pre-trained language model by initially freezing most layers and training only task-specific components, emphasizing the importance of controlled freezing to retain general linguistic knowledge captured during pre-training.
3.3 Formally verifying the fairness of language models
To evaluate the robustness and fairness of transformer-based models, we use formal verification based on zonotopes (Appendix A). Given a sentence, we want to verify that the prediction of the model stays the same even if the words are replaced by synonyms or gender-related terms. Please note that a brute-force technique quickly becomes computationally infeasible due to combinatorial explosions kumar2024certifying , particularly for longer sentences and words with many synonyms.
@@ 3.3 Formally verifying the fairness of language models
Thus, we capture synonyms (or gender-related terms) by modeling $\ell_{\infty}$-balls around each word representation in the embedding space as zonotopes. If these perturbation balls are large enough, they capture the synonyms as we pre-trained the embedding space such that these are in close proximity (Section 3.1). If a model is fair, we should be able to verify that the model does not change its prediction under large enough perturbations.
In particular, we perturb the embedding of each word in a sentence by some perturbation radius $\epsilon>0$. We then employ binary search to find the largest $\epsilon_{\text{max}}$ for which we can still verify that the prediction remains unchanged. Once $\epsilon_{\text{max}}$ is obtained, we assess if all synonyms of the words of interest lie within the computed perturbation ball. This can be easily checked by determining a threshold $\delta$ given by the largest $\ell_{\infty}$ distance between each word $i$ and its respective synonyms in the embedding space. Let $\hat{z}_{i}$ be the embedding of word $i$, then
@@
This verification setup returns a formal fairness certificate: Specifically, we construct input perturbation sets that include all fairness-sensitive terms, including different gender terms or different paraphrasing of toxic languages, and verify that no variation or substitution within the entire gender or toxic word set can alter the prediction of the model. The fairness is certified when the output of the model remains invariant across all perturbations.
@@
We can assess the overall fairness of a model by checking if such a certificate can be obtained for all sentences in our dataset. In particular, let there be $N$ sentences in the dataset, then the fairness score $\Psi$ can be computed by
@@
where $\mathbf{1}_{(\cdot)}$ is an indicator if the certificate is obtained for sentence $l$. Trivially, the closer $\Psi$ is to $1$, the fairer is a model in its predictions.
@@ 4.1 Datasets
To mitigate gender bias in the learned representations, we employ transfer learning to pre-train a Keras embedding layer using a curated set of gender-specific words bolukbasi2016man . The model is trained to cluster gendered word pairs (e.g., manwoman, malefemale, kingqueen) into a shared semantic class, minimizing the distance between them in the embedding space, while ensuring that unrelated words are separated. This strategy helps the model internalize semantic neutrality with respect to gender, which is critical for downstream fairness-sensitive tasks.
The dataset used in gender bias mitigation transformer training in this study is the Adult Census dataset adult_2 , including demographic information from the 1994 U.S. Census. The task is to predict whether an individual’s income exceeds \$50,000 per year, revealing potential biases in predictions based on race, gender, and other sensitive features. This dataset has been widely used in the context of fairness in machine learning lahoti2020fairness ; friedler2019comparative ; patel2024design ; le2022survey . Detailed preprocessing steps on the dataset can be found in Appendix B.
@@
The prevalence of toxic comments on social networking sites poses a significant threat to the psychological well-being of online users, which is why moderation is crucial pavlopoulos2020toxicity ; martens2015toxicity . To address this challenge, researchers have turned to machine learning algorithms as a means of categorizing and identifying toxic contents patel2025detecting ; 10.1145/3038912.3052591 .
The pre-training of embedding layers follows the same framework as for the gender bias mitigation experiment, using bad word collections from Google’s publicly available bad words list [^2] https://code.google.com/archive/p/badwordslist/downloads. For the training and evaluation of the transformer-based model, the Jigsaw Toxic Comment Classification dataset jigsaw2017toxic was utilized. This dataset has been widely adopted in various studies patel2025detecting ; chakrabarty2020machine ; revelo2023classification ; marchiori2022bias ; gupta2022machine . In this study, we use a curated version of the Jigsaw toxic comments dataset hateoffensive designed to evaluate the severity of toxic language by incorporating annotator agreement probabilities across three categories: neutral, offensive, and hate speech. We formulated a binary classification task, where neutral comments are labeled as non-toxic (label = 0), and both offensive and hate speech comments are labeled as toxic (label = 1).

### v2

@@ 1 Introduction
As large language models become integral to academia and industry, particularly in high-stakes decision-making and safety-critical domains, ensuring their robustness, and trustworthiness is crucial (García, 2023; Huang et al., 2024). While their transformer-based architecture with attention mechanisms has driven widespread adoption (Vaswani et al., 2017), concerns about robustness and safety persist, especially in tasks like fairness enforcement and content moderation.
A key challenge is their susceptibility to adversarial perturbations (Goodfellow et al., 2015b; Szegedy et al., 2014), where small perturbations, such as synonym substitutions or paraphrasing, can manipulate model behavior (Alzantot et al., 2018; Behjati et al., 2019; Li et al., 2019; Liang et al., 2018; Jin et al., 2020; Hsieh et al., 2019). While such perturbations may be adversarially crafted in settings like toxicity detection (Bespalov et al., 2023; Kruschwitz and Schmidhuber, 2024), they can also arise naturally in fairness-critical tasks such as salary prediction, where gender bias may occur, highlighting the need for robust gender bias mitigation (Tumlin et al., 2024). Many natural language processing tasks operate in discrete input spaces (language), where verifying robustness by enumerating all synonym substitutions becomes computationally infeasible due to combinatorial explosions. This holds especially for longer text sequences and words with many synonyms, which makes adversarial robustness verification particularly challenging (Bonaert et al., 2021; Huang et al., 2019). Specifically, the number of possible 20-token-long adversarial sequences drawn from a 32k-token (BERT-size) vocabulary is equal to $32\text{k}^{20}\approx 10^{90}$, a figure that exceeds the estimated number of atoms in the observable universe (Kumar et al., 2024).
@@ 2 Background
Figure 1: Overview of the framework explained in Section 3 involving pre-training (step 1 and 2), model training (step 3), and formal verification (step 4). The transformer architecture in step 3 is from (Vaswani et al., 2017).
We consider a general neural network classifier architecture that supports any element-wise nonlinear activation function. Given an input $x\in\mathbb{R}^{n}$, the classifier is represented by $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{c}$. A variety of tools have been developed to formally verify neural network properties, with adversarial robustness being the most extensively studied aspect (Brix et al., 2023b). The verification problem for neural networks can be formally stated as follows:
@@
Exact verification of robustness properties in neural networks is often computationally infeasible (Katz et al., 2017). Therefore, sound but incomplete methods are usually used for larger networks (Katz et al., 2019; Zhang et al., 2018; Bak, 2021; Lopez et al., 2023; Singh et al., 2018; Duong et al., 2024). In particular, we use zonotopes (Girard, 2005; Bonaert et al., 2021) to enclose the output of a neural network in this work, which has been shown to be a suitable foundation for analyzing larger neural networks (Singh et al., 2018; Gehr et al., 2018a; Mirman et al., 2018). As zonotopes offer an efficient and scalable way to represent perturbed input (or embedding) spaces, they have also been used to verify transformer-based models (Bonaert et al., 2021). A detailed explanation of zonotope-based neural network verification can be found in Appendix A. However, our work is agnostic to the underlying verification engine and others can be used as well (Brix et al., 2023a, 2024).
@@ 3 Methodology
This section presents the fairness verification framework in this work. As a running example, we illustrate the process using gender fairness verification in a salary prediction task. An overview of the framework is provided in Figure 1.
@@
We begin by pre-training an embedding layer on curated gendered word pairs, encouraging proximity between gendered pairs while maximizing distance from gender-neutral words. This enforces semantic neutrality and fairness at the embedding level through $\ell_{2}$ distance.
@@
The pre-trained embedding layer is integrated into a transformer-based language model and kept frozen during downstream training of the model, and then the language model is trained on a fairness-critical task involving gender bias mitigation.
@@ 3.1 Pre-training the embedding model
Since many AI-integrated automated decisionsication) can significantly impact people’s lives, it is paramou (including which individuals will receive jobs, loans, and mednt to assess and improve the ethics of the decisions made by these automated systems (Pessach and Shmueli, 2022; Mehrabi et al., 2021). A straightforward approach to promoting fairness in machine learning, defined as the absence of bias and discrimination, is to preclude the use of sensitive attributes during model training and inference (Lepri et al., 2018; Calders and Verwer, 2010; Kamiran et al., 2010; Barocas and Selbst, 2016). For instance, in designing a race-blind or gender-blind decision-making system, one might either exclude sensitive attributes, such as gender, from the feature set entirely (Lepri et al., 2018), or alternatively, represent them with similar embeddings that capture shared semantic meaning, placing them in the same semantic class to ensure equivalent treatment by the model (Bolukbasi et al., 2016). The corresponding experiments for both these scenarios are presented in Section 4.2.
@@
Word embedding models map words to vectors based on the assumption that words appearing in similar contexts share similar meanings (Pessach and Shmueli, 2022; Mikolov et al., 2013a, b). They capture semantic relationships and are widely used in natural language processing tasks (Pessach and Shmueli, 2022). However, studies have shown that embeddings can encode biases (Bolukbasi et al., 2016; Brunet et al., 2019; Caliskan et al., 2017b; Zhao et al., 2018); for instance, computer programmer may appear closer to male than female, raising concerns about potential discrimination in AI-based decision-making systems.
@@
This work employs pre-training-based transfer learning to train an embedding layer from scratch and cluster gender-related words while distancing unrelated ones. This offers greater transparency, controllability, and verifiability, enabling direct mitigation of gender bias during embedding layer training by mapping all gender-related words to a cohesive region in the embedding space. This semantic alignment enables formal verification. Similarly, this framework can also be applied to adversarial substitutions in the context of toxicity detection, ensuring robust censorship of inappropriate content such as profanity, hate speech, and abusive or otherwise harmful language. The approach is motivated by the need to better capture semantic relationships—especially involving gender, synonyms, or toxic terms—that are often missed by backpropagation alone [^1] Transfer learning is necessary because some embedding layers (e.g., Keras) are randomly initialized (Chollet and contributors, 2023), and training alone often fails to capture complex relationships like gender or toxicity.. Prior work supports this, showing embeddings trained on task-specific data can outperform static embeddings like Word2Vec (Bogale Gereme and Zhu, 2021). Similar practices are seen in domains such as fake news detection (Liu and Wu, 2018), highlighting the value of transfer learning to improve task performance and propagate semantic knowledge (Zhuang et al., 2021).
@@
The effectiveness of machine learning models relies heavily on data representation, with certain feature sets presumed tocapture core characteristics across tasks (Le-Khac et al., 2020). A good representation promotes intra-class compactness and inter-class separability (Khosla et al., 2020; Jaiswal et al., 2020; HaoChen et al., 2021; Wang and Liu, 2021). Contrastive learning has proven highly effective in enhancing representations across NLP, visual computing, and beyond (Jaiswal et al., 2020; Caron et al., 2020; He et al., 2020; Chen et al., 2020a; HaoChen et al., 2021; Saunshi et al., 2019; Chen et al., 2020b, e; He et al., 2020; Chopra et al., 2005). To learn semantically meaningful embeddings, we employ a contrastive loss that separates similar and dissimilar word pairs in the embedding space.
@@
The loss function used in this work is inspired by the formulation introduced in (Chopra et al., 2005) and has been adapted to suit our specific task. Here, the predefined margin $m>0$ specifies the minimum required separation between dissimilar pairs, while the hyperparameter $\alpha\geq 0$ regulates the influence of an additional linear attraction term applied to similar pairs. The loss encourages the model to minimize distances between similar pairs while enforcing a separation of at least $m$ units between dissimilar pairs in the embedding space. Examples of gender-related terms are listed in Section B.1, with contrastive loss details in Section B.2 and word pair construction in Section B.3. Additional pre-training with synonym and antonym pairs uses WordNet’s Synset structure (Miller, 1995) to model semantic similarity and contrast.
@@ 3.2 Training the language model
The pre-trained embedding layer is subsequently integrated into a transformer-based language model, with the architecture proposed in (Vaswani et al., 2017) and kept frozen and not updated via backpropagation during training to preserve linguistic knowledge, retaining general linguistic knowledge captured during pre-training, following common practice in models such as ELMo and ULMFiT (Peters et al., 2018; Howard and Ruder, 2018).
3.3 Verifying the fairness of language models
To evaluate the robustness and fairness of models, we use formal verification based on zonotopes (Appendix A). Given a sentence, we want to verify that the prediction of the model stays the same even if the words are replaced by synonyms or gender-related terms. Please note that a brute-force technique quickly becomes computationally infeasible due to combinatorial explosions (Kumar et al., 2024), particularly for longer sentences and words with many synonyms.
@@ 3.3 Verifying the fairness of language models
Thus, we capture synonyms (or gender-related or equivalent toxic terms) by modeling $\ell_{\infty}$-balls around each word representation in the embedding space as zonotopes. If these perturbation balls are large enough, they capture the synonyms as we pre-trained the embedding space such that these are in close proximity (Section 3.1). A fair model should maintain consistent predictions under sufficiently large perturbations.
We perturb the embedding of each word in a sentence by some perturbation radius $\epsilon>0$. We then employ binary search to find the largest $\epsilon_{\text{max}}$ for which we can still verify that the prediction remains unchanged. Once $\epsilon_{\text{max}}$ is obtained, we assess if all synonyms of the words of interest lie within the computed perturbation ball. This can be easily checked by determining a threshold $\delta$ given by the largest $\ell_{\infty}$ distance between each word $i$ and its respective synonyms in the embedding space. Let $\hat{z}_{i}$ be the embedding of word $i$, then
@@
Algorithm 1 Formal Fairness Verification
0: Dataset $\{S_{1},\dots,S_{N}\}$, model $f$, embedding $\hat{z}$, synonym sets $\mathcal{N}(i)$ for each word $i$
0: Fairness score $\Psi\in[0,1]$
[…]
2:for $l=1$ to NN do
3:  $\epsilon_{\text{max},l}\leftarrow$ find max. verifiable perturbation
4:  $\delta_{l}\leftarrow\max_{i\in S_{l}}\max_{j\in\mathcal{N}(i)}\left\lVert\hat{z}(i)-\hat{z}(w_{j})\right\rVert_{\infty}$
5:  if $\delta_{l}\leq\epsilon_{\text{max},l}$ then
6:   $c\leftarrow c+1$
7:  end if
8:end for
[…]
10:return $\Psi$
This verification setup returns a formal fairness certificate: Specifically, we construct input perturbation sets that include all fairness-sensitive terms, including different gender terms or different paraphrasing of toxic language, and verify that no variation (of gender attributes) or substitution (of semantically equivalent toxic terms) within the entire perturbation set can alter the prediction of the model. Fairness is then certified if no such perturbation exists.
@@
We can assess the overall fairness of a model by checking the percentage of all of the sentences in our test dataset, such a certificate can be obtained. In particular, let there be $N$ sentences in a held-out test dataset (not seen during model training), then the fairness score $\Psi$ can be computed by
@@
where $\mathbf{1}_{(\cdot)}$ is an indicator if the certificate is obtained for sentence $l$. Trivially, A higher $\Psi$ (closer to 1) implies a fairer model. The whole verification procedure is shown in Alg. 1.
@@ 4.1 Datasets
To mitigate gender bias in the learned representations, we employ transfer learning to pre-train a Keras embedding layer using a curated set of gender-specific words curated in prior work (Bolukbasi et al., 2016). The model is trained to cluster gendered word pairs (e.g., manwoman, malefemale, kingqueen) into a shared semantic class, minimizing the distance between them in the embedding space, while ensuring that unrelated words are separated. This strategy helps the model internalize semantic neutrality with respect to gender, which is critical for downstream fairness-sensitive tasks. Detailed preprocessing steps on the pre-training dataset can be found in Appendix B.
The dataset used in the model training for the task of gender bias mitigation in this study is the Adult Census dataset (Becker and Kohavi, 1996), including demographic information from the 1994 U.S. Census. The task is to predict whether an individual’s income exceeds \$50,000 per year, revealing potential biases in predictions based on race, gender, and other sensitive features. This dataset has been widely used in the context of fairness in machine learning (Lahoti et al., 2020; Friedler et al., 2019; Patel and Jadhav, 2024; Le Quy et al., 2022). Detailed preprocessing steps, including the serialization of tabular data by converting to a sentence dataset so that it is consumable by transformers (Badaro et al., 2023), with prior literature supporting this method (Chen et al., 2020d; Suadaa et al., 2021; Chen et al., 2020c), can be found in Section B.4.
@@
The prevalence of toxic comments on social networking sites poses a significant threat to the psychological well-being of online users, which is why moderation is crucial (Pavlopoulos et al., 2020; Märtens et al., 2015). To address this challenge, researchers have turned to machine learning algorithms as a means of categorizing and identifying toxic contents (Patel et al., 2025; Wulczyn et al., 2017).
The pre-training of embedding layers follows the same framework as that of the gender bias mitigation experiment, using a bad words collection from Google’s publicly available bad words list [^2] https://code.google.com/archive/p/badwordslist/downloads. For the training and evaluation of the language model, the Jigsaw Toxic Comment Classification dataset (Adams et al., 2017) was utilized. This dataset has been widely adopted in various studies (Patel et al., 2025; Chakrabarty, 2020; Revelo-Bautista et al., 2023; Marchiori Maneba et al., 2022; Gupta et al., 2022). In this study, we use a curated version of the Jigsaw toxic comments dataset (Davidson et al., 2017) designed to evaluate the severity of toxic language by incorporating annotator agreement probabilities across three categories: neutral, offensive, and hate speech. We formulated a binary classification task, where neutral comments are labeled as non-toxic (label = 0), and both offensive and hate speech comments are labeled as toxic (label = 1).